Mission Operations Days: 2001/265 to 2001/271
September 22, 2001 00:00:00 GMT to September 28, 2001 00:00:00 GMT
Terra Spacecraft and MODIS Instrument Status:
|Terra (AM-1) is in Normal Ops Mode|
MODIS is in Science Mode - A Side
|Blackbody||A On (290 K); B Off||Nominal|
|Calibration Electronics||A On; B Off||Nominal|
|Control Processor||A On; B Off||Nominal|
|Nadir Door||Unlatched, Open||Nominal|
|Space View Door||Unlatched, Open||Nominal|
|Solar Diffuser Door||Unlatched, Closed||Nominal|
|FDDI Formatter||A On; B Off||Nominal|
|FDDI Port||A On; B Off||Nominal|
|FIFO Memory||#1 & #2 On; #3 & #4 Off||Nominal|
|Format Processor||A On; B Off||Nominal|
|Power Supplies||#1 On; #2 Off||Nominal|
|PV VIS FPA||A On; B Off||Nominal|
|PV NIR FPA||A On; B Off||Nominal|
|PV SMIR FPA||A On; B Off||Nominal|
|PV LWIR FPA||A On; B Off||Nominal|
|PC LWIR FPA||A On; B Off||Nominal|
|RC Outgas Heaters||All Off (CS, IS, OS)||Nominal|
|RC LWIR FPA Heater||On (83 K)||Nominal|
|RC SMIR FPA Heater||Off||Nominal|
|Scan Assembly||A On; B Off||Nominal|
|SDSM||A Off; B Off||Nominal|
|SRCA||A Off; B Off||Nominal|
|PS1 Survival Heater||Enabled||Nominal|
|PS2 Survival Heater||Enabled||Nominal|
|Timing Generator||A Off; B On||Nominal|
|Flight Software||Rev BD + 2 set of patches||Nominal|
|Inhibit Ids Set||None||Nominal|
|TMON 66 (SRCA A)||Disabled||Nominal|
|TMON 67 (SRCA B)||Disabled||Nominal|
|TMON 68 (FIFO)||Enabled||Nominal|
|TMON 69 (EPS Current)||Disabled||Nominal|
This Week's Completed MODIS Activities:
Saturday, September 22, 2001
Sunday, September 23, 2001
Monday, September 24, 2001
Tuesday, September 25, 2001
2001/268 21:17:38 Real-Time Executes MOD_ALARMS_CP_ST_06 to clear formatter error address flag. Flag is at hex A (10).
Wednesday, September 26, 2001
2001/269 20:47:46 ATC OA16: SDSM SCREEN
2001/269 22:26:39 ATC OA15: SDSM OPEN
Thursday, September 27, 2001
2001/270 20:08:33 – 20:09:55 Real-Time Dumps memory addresses for CP_ST_06 formatter errors.
Friday, September 28, 2001
This Week's Scheduled MODIS Activities Not Completed:
Upcoming MODIS Events:
Friday, September 28, 2001
Saturday, September 29, 2001
Sunday, September 30, 2001
Monday, October 1, 2001
Tuesday, October 2, 2001
Wednesday, October 3, 2001
Thursday, October 4, 2001
Friday, October 5, 2001
There will be a drag makeup maneuver on day 275/185100. Duration of burn will be 27 seconds.
The next Lunar Calibration roll maneuver is scheduled for day 2001/280 (October 7, 2001) with data collection from 07:10:54 to 07:17:24. The roll angle will be -11.3321 degrees.
Formatter Error Events:
The FR error counter (MOD_SS_CP_STATUS_08) has “rolled over” 9 times since the A-Side restart following the PS-2 Anomaly, and 7 times since the latest string of errors started on August 15. This is a 16-bit signed counter, therefore it “rolls over” every 32,768 counts. At present (2001/271-16:10:00), the counter is 614, which means that 295,526 errors have been seen since the A-Side restart, and 215,718 since August 15.
A cyclic, increasing rate of Formatter errors has been detected. The increase in Formatter error rates occur at the same times we see decreases in the MOD_TP_FR_A_ENGINE, MOD_TP_ME_NX_HTSINK, and MOD_TP_ME_CHAS_TOP temperatures. The current trend of these mnemonics is a gradual decrease in temperature.
During the recent SRCA activities, the rate of FR errors was significantly reduced. At the same time, the average temperatures of the three telemetry points listed above each rose approximately 1 degree C.
MS Excel spreadsheets are available for download from:
These files contain the times for each FR error count as well as the three temperatures (from day 240 on) mentioned above. This file will be periodically updated with the most recent telemetry sets available. In this file, you can see a graph of this cyclic, increasing error rate. If you have any special requests, please let us know.
The current of FR errors is now approaching 20,000 a day. As of yet, no problems with the science data have been reported. If the error rate continues to increase, there is the possibility of brief interruptions in science data as the FR attempts to deal with the large amount of errors. At present we have no indication to what this threshold level for science data loss may be.
We are still in a "wait and see" mode right now. However, everyone should be encouraged to consider options (i.e. B-side) before we have no choices left.
General Instrument Comments:
MODIS is in Science Mode on the A-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).
Note: SD/SDSM calibrations were planned to be reduced to once per two weeks. This plan has been revised; SD/SDSM calibrations will continue to be scheduled once each week until the end of November.
MODIS Telemetry Trends:
Trending has resumed with the instrument on the A-side. With the currently decreasing beta angle, the amount of focal plane heater margin is increasing. The intermediate stage cooler temperature is stable.
The increase in formatter error events is also being trended (as described above) to aid in any future operational decisions necessary to maintain the health of the instrument and its science data.
Non-MODIS Significant Events:
Limited Life Item Status:
SRCA 10W Lamp #1: 189.7 of 500 hours
SRCA 10W Lamp #2: 145.3 of 500 hours
SRCA 10W Lamp #3: 154.5 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours
SRCA 1W Lamp #1: 558.2 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours
Solar Diffuser Door: 1715 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements